40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jul 2 11:21:24 2013, Erica, DailyProgress, fiber optic, mode matching snip_of_beam_profiling_6_27.JPGP1020518.JPG
    Reply  Wed Jul 3 09:46:48 2013, Erica, DailyProgress, fiber optic, mode matching 130703-095332.jpg
       Reply  Thu Jul 11 01:40:19 2013, Erica, DailyProgress, fiber optic, mode matching P1020690.JPGP1020692.JPG
          Reply  Fri Jul 12 00:51:39 2013, Erica, Notes, fiber optic, mode matching P1020705.JPGP1020703.JPG7_11_13_beam_profile.PNG7_11_13_beam_from_fixed_lens.PNG
             Reply  Tue Jul 16 10:27:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 
                Reply  Tue Jul 16 11:02:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 130805-120628.jpg
                   Reply  Wed Jul 17 11:58:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 7_16_13_determining_polarization_axis.PNGTEK0000.PNG
                      Reply  Wed Jul 17 18:29:37 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                         Reply  Sat Jul 20 23:39:09 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                            Reply  Tue Jul 23 10:29:12 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam shielded_10s_TEK0005.PNGunshielded_10s_TEK0027.PNG
                               Reply  Tue Jul 23 11:43:00 2013, Evan, DailyProgress, Electronics Equipment, Zach's op-amp recommendations 
                               Reply  Fri Jul 26 11:07:44 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table P1030056.JPGP1030057.JPGP1030059.JPGP1030056.JPG
                                  Reply  Fri Jul 26 11:22:45 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                     Reply  Fri Jul 26 11:45:28 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                        Reply  Mon Jul 29 13:30:26 2013, Erica, DailyProgress, BEAT, adding resistor, capacitor, and sockets w_o_capacitor_7_26_13.JPGcapacitor_7_26_13.JPGP1030066.JPGP1030070.JPG
                                           Reply  Tue Jul 30 12:23:41 2013, Erica, DailyProgress, BEAT, addressed noise problem, DC signal 
                                              Reply  Thu Aug 1 00:25:02 2013, Erica, DailyProgress, fiber optic, insulating foam 731_7graph.PNG731_11graph.PNGP1030318.JPG
                                                 Reply  Fri Aug 2 12:38:41 2013, tara, DailyProgress, fiber optic, insulating foam chmbrtmp.JPGfringe.JPG
                                                    Reply  Tue Aug 6 12:31:24 2013, Erica, DailyProgress, fiber optic, insulating foam P1030403.JPG
                                                       Reply  Wed Aug 7 13:27:00 2013, Erica, DailyProgress, fiber optic, more foam, calculations for convertion from voltage to frequency P1030409.JPGP1030417.JPGP1030420.JPG
                                                          Reply  Wed Aug 7 18:53:20 2013, Erica, DailyProgress, fiber optic, noise budget, Matlab notes. noise_budget.PNGP1030427.JPG
                                                             Reply  Tue Aug 13 12:39:36 2013, Erica, DailyProgress, fiber optic, rewrote calculations for converting from voltage noise to frequency noise P1030458.JPGP1030459.JPGP1030460.JPG
                                                       Reply  Tue Aug 13 12:28:20 2013, Erica, Notes, Electronics Equipment, new AD 590 circuit and its noise calculation, beam alignment P1030461.JPGP1030454.JPGP1030456.JPGP1030457.JPGAD590circuit2.PNG
                                                          Reply  Fri Aug 16 10:35:44 2013, Erica, Notes, Electronics Equipment, AD590 msrmt, conversion to freq noise 
                   Reply  Tue Jul 30 16:19:32 2013, Erica, Notes, fiber optic, setup for beam recombination new_setup.PNGbeam_recombination2.PNG
                   Reply  Mon Aug 5 12:21:00 2013, Erica, Notes, fiber optic, calculation: phase change for fiber 130805-121449.jpg
    Reply  Tue Jul 30 15:48:47 2013, Erica, DailyProgress, fiber optic, noise for GYRO 
Message ID: 1224     Entry time: Wed Jul 3 09:46:48 2013     In reply to: 1222     Reply to this: 1235
Author: Erica 
Type: DailyProgress 
Category: fiber optic 
Subject: mode matching 
July 2, 2013

Started writing up the progress report due July 9. Need to find more papers about the different types of noise from a fiber.


I talked to Eric Q about the crackle lab. In the interferometer at LIGO, there are blade springs with mirrors that hang from them. The laser reflects off of these mirrors; a horizontal displacement of the mirror leads to a vertical displacement of the blade spring. Noise in the laser can lead to perturbations of the mirror, and therefore, perturbations of the blade springs, which can mask real signals. His setup is a Michelson interferometer with the mirrors placed at the end of masses hanging from blade springs. See attachment 1 for a diagram of the setup. His requirement for noise is 10^-15 m/sqrt(Hz) for the spring, with a displacement of about 1 cm, operating at a frequency of about 100 Hz. Using the equation in attachment, it says we need a frequency noise of 10^(-11) Hz/sqrt(Hz) which is much lower than what I expected so I will have to ask again to clarify.

Recombined the beams from the fiber and the laser and set up a photodiode. Placed the collimator from the fiber optic on a translational stage because it allows us to minutely change the length in order to more easily change the relative phase difference between the two beams. Talked to Evan about finding the beat frequency. We want to adjust the lengths of the two beams such that the phase between them is pi/2, so there is completely destructive interference => wL/c = (2n+1)pi/2.
Attachment 1: 130703-095332.jpg  123 kB  Uploaded Wed Jul 3 10:59:51 2013  | Hide | Hide all
130703-095332.jpg
ELOG V3.1.3-